Green Caterpillar Garden Supplies Inc. is a hypothetical company. Suppose it has the following balance sheet items reported at the end of its first year of operation. For the second year, some parts are still incomplete. Use the information given to complete the balance sheet Green Caterpillar Garden Supplies Inc.Balance Sheet for Year Ending December 31 (Millions of Dollars) Year 2 Year 1 Year 2 Year 1 Assets Current assets: Cash and equivalents Accounts receivable Inventories Total current assets Net fixed assets: Net plant and equipment Liabilities and equity Current liabilities: Accounts payable Accruals Notes payable Total current liabilities Long-term debt 2,953 8,662 $19,687 $6,457 2,363 6,930 $15,750 $0 410 2,324 $0 0 2,187 $%2,187 8,203 6,563 $19,250 Tl debt $10,937 $8,750 Common equity Common stock Retained earnings Total common equity Total liabilities and equity 21,328 17,063 9,187 $26,250 $35,000 $32,813 Total assets $43,750 $35,000 $43,750Given the information in the preceding balance sheet-and assuming that Green Caterpillar Garden Supplies Inc. has 50 million shares of common stock outstanding-read each of the following statements, then identify the selection that best interprets the information conveyed by the balance sheet. Statement # 1: Green Caterpillars accumulated owed financial obligations decreased from Year 1 to Year 2. This statement is because: Accruals actually increased from $0 in Year 1 to $410 million at the end of Year 2 The accounts payable account remained $0 for both Years 1 and 2 Long-term debt decreased from $2,324 million at the end of Year 1 to $2,187 million by the end of Year 2Statement #2: On December 31 of Year 2, Green Caterpillar Garden Supplies Inc. had $8,072 million of actual money that it could have spent immediately This statement is ,because: Green Caterpillars Year 2 cash and equivalents balance is $20,317 The funds recorded in Green Caterpillars accounts receivable account represents funds that are either cash or can be converted into cash almost immediately The funds recorded in Green Caterpillars cash and equivalents account represents funds that are either cash or can be converted into cash almost immediatelyStatement #3: If Green Caterpillar ever goes bankrupt, its common stockholders will be paid off first, then its debtholders and preferred stockholders This statement is ,because: Common shareholders are treated as residual investors Debtholders and preferred shareholders are considered residual investors Debtholders are treated as residual investors Based on your understanding of the different items reported on the balance sheet and the information they provide, if everything else remains the same, then the cash and equivalents item on the current balance sheet is likely to common stock if the firm issues $3 million of new
Answer #1

1: Cash and cash equivalents = Total current assets – AR- Inventories

= 19687-2953-8662

=8072

2: Net plant and equipment = Total assets- current assets

= 43750-19687

=24063

3: Total current liabilities = Accounts payable + Accruals + notes payables

= 0+410+2324

=2734

4: Retained earnings = Total common equity – common stock

= 32813- 21328

=11485

5: Statement 1 is FALSE because

Accruals actually increased from year 1 to year 2

(Accruals have increased. Current liabilities have also increased )

6: Statement 2 is TRUE because

Option 3 is the answer

(The cash and cash equivalents balance in year 2 is 8072)

7: Statement 3 is FALSE because

Option 1 is true

(Equity shareholders are paid off in the end)

8: Increase

(If new stock is issued, there will be cash inflow)